### v1

# Unveiling the Potential of Multimodal Retrieval Augmented Generation with Planning

Abstract.
Multimodal Retrieval Augmented Generation (MRAG) systems, while promising for enhancing Multimodal Large Language Models (MLLMs), often rely on rigid, single-step retrieval methods. This limitation hinders their ability to effectively address real-world scenarios that demand adaptive information acquisition and query refinement. To overcome this, we introduce the novel task of Multimodal Retrieval Augmented Generation Planning (MRAG Planning), focusing on optimizing MLLM performance while minimizing computational overhead. We present CogPlanner, a versatile framework inspired by human cognitive processes. CogPlanner iteratively refines queries and selects retrieval strategies, enabling both parallel and sequential modeling approaches. To rigorously evaluate MRAG Planning, we introduce CogBench, a new benchmark specifically designed for this task. CogBench facilitates the integration of lightweight CogPlanner with resource-efficient MLLMs. Our experimental findings demonstrate that CogPlanner surpasses existing MRAG baselines, achieving significant improvements in both accuracy and efficiency with minimal computational overhead.
Multimodal Retrieval Augmented Retrieval, Query Planning, Multimodal Large Language Model, Visual Question Answering
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Figure 1. An example of current MRAG system with the benefits of incorporating MRAG Planning.
Retrieval Augmented Generation (RAG) has proven to be a significant advancement in enhancing the generation quality of large language models (LLMs) (Izacard and Grave, 2020; Lewis et al., 2020). While RAG frameworks have primarily focused on textual information, the burgeoning need for more sophisticated systems demands the incorporation of multimodal data, including images and videos (Riedler and Langer, 2024). This evolution has given rise to Multimodal Retrieval Augmented Generation (MRAG), which aims to enable multimodal large language models (MLLMs) to access and leverage multimodal external knowledge sources, thereby mitigating hallucinations and enhancing the system reliability (Zhao et al., 2023a). However, challenges lie in the profound intricacies of real-world human interactions where the users tend to provide low-resolution screenshots (Menges et al., 2018) and succinct queries with heavy reliance on anaphoric references (Spink et al., 2001). This behavior necessitates accurate information-seeking and integration mechanisms to construct comprehensive responses. Existing MRAG frameworks generally adhere to a rigid pipeline characterized by a predetermined retrieval action, either exclusively textual or exclusively visual (Chen et al., 2022; Zhu et al., 2024), which manifests several critical limitations:
  • Blind Information Acquisition: As demonstrated in recent studies (Shi et al., 2023; Asai et al., 2023), the compulsive retrieval mechanism without proper consideration of necessity or relevance can introduce irrelevant contextual information that undermines the MLLM’s capability for accurate responses. Moreover, it neglects the inherent capabilities of MLLMs to reason and process multimodal data, rendering the retrieval step redundant or counterproductive.
  • Inadequate Query Formation: The visual incompleteness, textual ambiguity, and conciseness create a fundamental impediment that fails to retrieve pertinent information. Furthermore, the current single-step MRAG methodologies prove inadequate for addressing multi-hop queries that require multi-step reasoning process (Kil et al., 2024).
In response to these real-world limitations, we introduce a novel task, termed Multimodal Retrieval Augmented Generation Planning (MRAG Planning), which determines an optimal procedure for multimodal query processing, enabling judicious information seeking and integration. The objective is to maximize the capacity of the underlying MLLMs to generate a comprehensive and accurate response while minimizing the computational cost overhead to the MRAG system. To this end, the MRAG Planning task is formally decomposed into two interrelated sub-tasks: (1) Information acquisition which discerns the truly necessary information required by the MLLMs and devises the appropriate retrieval strategies accordingly. (2) Query reformulation which involves decomposing complex, multi-hop queries into manageable atomic sub-queries, and refining them for clarity and informativeness. Figure 1 illustrates a multi-hop reasoning example where the first step is to identify the game name AstroBot through an image search. The initial query is then decomposed into two sub-queries, each focusing on the sales data for the two respective games. Subsequent web searches are conducted to retrieve sales data and generate the final response. This example underscores the crucial need for dynamic planning procedures in MRAG systems where both order and selection of retrieval methods, and query reformulation are determined tailored to the specific characteristics of the multimodal query.
To address the challenges of MRAG Planning, we propose a novel framework, CogPlanner  [^1] We name our framework as CogPlanner because of the inherent cognitive process of humans., inspired by human cognitive processes. Just as humans synthesize and gather multimodal information to address complex queries, adapting to their cognitive abilities and knowledge base, CogPlanner emulates this behavior by dynamically determining a planning procedure in coordination with downstream MLLMs. CogPlanner operates through two core actions: query reformulation and retrieval action selection. Query reformulation involves breaking down a complex query into related sub-queries or refining the queries. The retrieval strategy, in turn, encompasses image search, text search, or none. When sufficient information is gathered, the framework refrains from further retrieval, culminating in the generation of a final response. This iterative procedure reflects an adaptive chain of actions tailored to the specific needs of each multimodal query. For instance, the optimal planning strategy in Figure 1 follows (Query Refinement, Image Search) $\rightarrow$ (Query Decomposition, Text Search). CogPlanner supports two distinct modeling approaches: parallel and sequential modeling, each differs in the order of the query reformulation and retrieval action selection actions.
In conjunction with the introduction of the CogPlanner framework, we present CogBench, a comprehensive dataset tailored to the decision-making processes central to the MRAG Planning task. CogBench consists of over 7,000 data samples, with a high-quality test set of 401 samples. The creation of CogBench involves two steps: (1) Query Collection, wherein we enrich with collected screenshots and multi-hop reasoning queries demanding long-form answers, and (2) MRAG Planning and Generation, wherein we apply CogPlanner with GPT-4o (Achiam et al., 2023) and human annotations. In addition to its essential role in evaluation, the development of CogBench facilitates the design of specialized fine-tuning strategies aimed at bolstering the decision-making capabilities of smaller, resource-efficient MLLMs. By utilizing the CogBench training set, we achieve lightweight integration of the Qwen2-VL-7B-Instruct (Wang et al., 2024) model as the planning expert in CogPlanner. This integration, referred to as Qwen2-7b-VL-Cog, maintains its resource-efficient characteristics while enabling effective performance within the MRAG Planning context. In summary, the contributions of this work are as follows:
  • We thoroughly examine the limitations of current MRAG frameworks, specifically addressing the challenges of information acquisition and query reformulation. Building upon this, we formally define the task of Multimodal Retrieval Augmented Generation Planning (MRAG Planning), laying the groundwork for further research.
  • We introduce CogPlanner, a flexible, plug-and-play framework that incorporates two distinct modeling approaches, parallel modeling and sequential modeling.
  • We develop CogBench benchmark, tailored to the MRAG Planning task. It supports performance evaluation and facilitates finetuning strategies to enable the lightweight integration of resource-efficient MLLMs with CogPlanner. Experimental results demonstrate that CogPlanner achieves more than 15% improvements over various MRAG approaches while incurring less than 10% additional costs with Qwen-7b-VL.
2. Related Work
2.1. Query Processing in IR
Query processing is a critical aspect of Information Retrieval (IR) systems, directly influencing the efficiency and effectiveness with which relevant information is retrieved in response to user queries. Early IR systems rely on complex, multi-stage query processing pipelines, which incorporate a range of techniques, including query rewriting (Ma et al., 2023), intention detection (Arora et al., 2024), sentiment analysis (Zhang et al., 2023), and query expansion (Wang et al., 2023), among others. These pipelines often utilize human-defined heuristics to refine the query, enabling more precise document retrieval (Li et al., 2022; Lee et al., 2020). However, the advent of large language models (LLMs) has significantly transformed this approach. The exceptional expressive power and reasoning capabilities of LLMs allow them to effectively perform several traditional query processing tasks within a single, well-crafted prompt. This shift has been particularly notable in the context of RAG systems, where the primary challenge now lies in determining the most effective strategy for processing user queries.
2.2. Multimodal Retrieval Augmented Generation
RAG frameworks have demonstrated considerable success in various real-world applications (Li et al., 2022). However, their reliance on textual information presents a significant limitation, as it precludes the incorporation of crucial knowledge embedded within other modalities, such as images and videos. Multimodal Retrieval Augmented Generation (MRAG) seeks to address this limitation by equipping MLLMs with access to a broader spectrum of knowledge, encompassing up-to-date and domain-specific information (Zhao et al., 2023b). Empirical studies consistently demonstrate the effectiveness of MRAG systems (Zhao et al., 2023a) across various visual question answering (VQA) benchmarks (Wu et al., 2017). Recent advancements in MRAG have demonstrated notable progress. For instance, MuRAG (Chen et al., 2022) highlights how incorporating visual information retrieved from external sources significantly improves the system performance. Other prominent approaches include Plug-and-Play (Tiong et al., 2022), which transforms visual content into textual descriptions to facilitate integration with conventional text-based question-answering mechanisms. Additionally, RAMM (Yuan et al., 2023) enhances the generation process by incorporating both text-to-image retrieval and subsequent fusion of the representations for more accurate answer generation. Further innovations include Wiki-llava (Caffagni et al., 2024) and mR2AG (Zhang et al., 2024b), which enable retrieval from online knowledge bases, such as Wikipedia, via image-based queries, to provide more contextually informed responses to user queries. M2RAG (Ma et al., 2024) extends these efforts by enabling concurrent retrieval of both textual and visual elements in response to multimodal queries, allowing for more robust query understanding and generation capabilities. Additionally, benchmarks such as MRAG-Bench (Hu et al., 2024) and MMSearch (Jiang et al., 2024b) have been introduced to evaluate MRAG performance, particularly in tasks with image-to-image retrieval, addressing challenges related to incomplete or insufficient image data.
The prevailing methodologies predominantly adhere to a rigid, single-modality search paradigm. However, in authentic user scenarios, knowledge acquisition can originate from diverse sources, contingent upon the specific query and the underlying domain. To this end, we introduce the novel task of MRAG Planning. The primary goal of MRAG Planning is to systematically determine the most effective query processing strategy tailored to each multimodal query and the underlying MLLM.
3. Task Formulation
Consider a multimodal query $\mathcal{Q}_{0}=(q,v)$ where $q$ represents the textual component and $v$ represents the visual component (e.g. an image). The objective of MRAG is to retrieve pertinent information from a document collection $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{n}\}$, and generate cogent responses. Drawing parallels with human information processing, we introduce the task of MRAG Planning that interfaces intimately with the retrieval tools and downstream MLLMs in the MRAG systems, restructuring their information gathering mechanism. We formalize the MRAG system environment as a tuple $(\mathcal{G},\mathcal{I})$. Here, $\mathcal{G}$ refers to the goal conditions of assembling sufficient information to generate comprehensive and accurate responses. $\mathcal{S}$ represents the state, capturing the current set of information available, which may encompass queries and any retrieved documents. The initial state corresponds to the input multimodal query, $\mathcal{I}=\mathcal{S}_{0}=\mathcal{Q}_{0}$. The MRAG Planning task can thus be framed as a state transition function $\mathcal{F}$ that progresses from $\mathcal{I}$ toward the goal state through a chain of decisions. Formally, this transition process is defined as follows:
(1) $\mathcal{F}:\mathcal{S}\times\mathcal{P}\rightarrow\mathcal{S},$
where $\mathcal{P}$ represents the available decision space.
3.1. Multimodal Retrieval Augmented Generation Planning
In line with the human cognitive architecture’s capacity for knowledge integration, we conceptualize the MRAG Planning task as a dual optimization problem comprising two sub-tasks: information acquisition and query reformulation.
3.1.1. Planning Procedure
The decision space can thus be decomposed as $\mathcal{P}=(\mathcal{A},\mathcal{Q})$, where $\mathcal{A}$ represents the information acquisition strategy and $\mathcal{Q}$ denotes the query reformulation result. To accommodate multi-hop reasoning queries, the planning process unfolds iteratively across $T$ rounds under the assumption of Markovian state transitions. Specifically, at each iteration $t$, the decision is determined by the current available information state:
(2) $\mathcal{S}_{t}\rightarrow(\mathcal{A}_{t},\mathcal{Q}_{t}).$
The subsequent state is updated naturally as follows:
(3) $\{\mathcal{S}_{t},(\mathcal{A}_{t},\mathcal{Q}_{t})\}\rightarrow\mathcal{S}_{t+1},$
where $\mathcal{S}_{t}$ may encompass the historical queries or the retrieved document elements, denoted as $\mathcal{D}^{t}$.
3.1.2. Information Aquisition
Recognizing the inherent limitations of MLLMs in terms of specific knowledge gaps, an information acquisition mechanism is imperative to supplement the MLLM’s knowledge base. We define $\mathcal{A}$ comprising three distinct retrieval operations, text search, image search, and non-search. The optimal retrieval action is determined based on the quality of the available multimodal information and the estimated utility of additional context from external knowledge sources:
(4) $\mathcal{A}_{t}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\mathcal{F}_{IA}(a|\mathcal{S}_{t}).$
Post action selection, we proceed with in-document retrieval, which identifies and extracts pertinent elements under the selected retrieval strategy, yielding retrieved document elements $\mathcal{D}^{t}$.
3.1.3. Query Reformalization
To address the ambiguity and potential incompleteness in queries, we refine the queries by leveraging both textual and visual cues within the multimodal query. For complex queries necessitating multi-hop reasoning, we employ a decomposition strategy that preserves semantic relationships while breaking down the query into manageable sub-queries. This process can be expressed as:
(5) $\mathcal{Q}_{t}=\mathcal{F}_{QR}(\mathcal{Q}_{t-1},\mathcal{S}_{t})=\{\mathcal{Q}_{t,1},\ldots,\mathcal{Q}_{t,N_{t}}\},$
where $N_{t}$ represents the cardinality of the decomposed and refined query set at iteration $t$.
3.1.4. Generation
Once the final state $\mathcal{S}_{T}$ is deemed sufficiently informed, the planning process culminates in response generation. The response is synthesized by incorporating the initial query, the final refined query, and the relevant document elements:
(6) $\text{Response}=\mathcal{F}_{Gen}(\mathcal{Q}_{0},\mathcal{Q}_{T},\mathcal{D}^{T}),$
where $\mathcal{F}_{Gen}$ represents the MLLM generator.
Figure 2. The overall framework of CogPlanner.
4. Methodology
4.1. Baselines
Existing MRAG methodologies predominantly rely on a fixed information acquisition pipeline, characterized by a single-modality retrieval action performed in a single turn. These methods can be broadly classified into two distinct categories:
  • Fixed textual retrieval: These methods closely resemble traditional RAG frameworks. They employ textual queries to retrieve relevant documents. Subsequently, the original user query and the retrieved information are utilized as complementary inputs to generate the final answer.
  • Fixed visual retrieval: These approaches prioritize visual information, leveraging visual queries to retrieve relevant images and their associated captions. The MLLM then integrates both the original user query and the augmented retrieved images to produce the final answer.
While these fixed workflows have been effective within certain domains, they exhibit limited flexibility. The rigid, modality-specific structure hinders their ability to adapt to other contexts, or the non-search scenario where fixed search may bring extra noise. It undermines their potential for broader applicability in dynamic, real-world scenarios. In our experimental settings, we utilize these two methodologies as baseline comparisons to evaluate the efficacy of our proposed CogPlanner.
4.2. CogPlanner
We propose CogPlanner, a flexible, plug-and-play framework that mirrors the human cognitive processes when handling complex multimodal queries. Our approach is motivated that this process exhibits both adaptability and efficiency in managing multimodal information acquisition and integration. When confronted with such complex multimodal queries, humans instinctively engage in a methodical, multi-step process that involves continuous evaluation of information requirements, determination of appropriate retrieval mechanisms for the missing knowledge, and decomposition and refinement of complex queries into manageable sub-components. What distinguishes this process is its iterative nature, where each step builds upon previous understanding, ultimately converging on an information-rich and query-clear state. Crucially, each individual naturally modulates their planning based on their knowledge base and cognitive capabilities.
CogPlanner operationalizes this cognitive architecture through an iterative decision-making framework. It dynamically orchestrates a chain of actions for each query, optimizing both effectiveness and efficiency in conjunction with the downstream MLLMs. As shown in Figure 2, the multimodal query — seeking up-to-date news about Luigi’s arrest — requires three processing rounds to gather sufficient information for the current GPT knowledge base. CogPlanner centers on two critical decisions in $\mathcal{P}$: (1) query reformulation, and (2) action selection among text search, image search, and non-search, corresponding to the sub-tasks outlined in Section 3.1. We implement the core state transition function $\mathcal{F}$ through a planning expert who makes these decisions at each iteration, completing the roles of $\mathcal{F}_{IA}$ and $\mathcal{F}_{QR}$. Then, the retrieval are invoked to assess and process relevant multimodal elements.
4.2.1. Planning Expert
We employ a MLLM as the planning expert. As formalized in Equation 2, at each iterative step, the expert analyzes the current multimodal content, evaluates the information gathered thus far, and subsequently determines the most appropriate follow-up retrieval action and reformulates the query. To accomplish this, we propose two distinct modeling paradigms - parallel modeling and sequential modeling - each differs in the order in which the decision-making occurs.
Parallel Modeling In the parallel modeling paradigm, the planning expert concurrently adjusts the query and determines the appropriate retrieval action. Specifically, it takes the current query alongside the information retrieved in the preceding iteration as inputs. The implementation employs two parallel threads of MLLM inference: one is responsible for query reformulation, while the other determines the next retrieval action. The primary advantage of this paradigm lies in efficiency. By enabling simultaneous decision-making, we achieve a streamlined decision chain and faster processing. This is particularly beneficial in real-world applications where response latency is a critical consideration. The parallel process can be formally expressed as:
(7) $\mathcal{Q}_{t}=$ $\mathcal{F}_{QR}(\mathcal{Q}_{t-1},\mathcal{D}^{t}),$
$\mathcal{A}_{t}=$ $\mathcal{F}_{IA}(\mathcal{Q}_{t-1},\mathcal{D}^{t}).$
Sequential Modeling The sequential paradigm, in contrast, implements an ordered two-step decision process. The first step involves query reformulation, wherein multimodal inputs, along with previously retrieved information, are leveraged to refine and decompose the query. These restructured queries then serve as the inputs for the subsequent stage, which entails an evaluation of the necessity of further retrieval actions. This paradigm facilitates a more nuanced understanding of information retrieval requirements, as the planning expert is presented with both the original query and its reformulated meta-queries. It mimics a reflective cognitive process where the planning expert jointly assesses whether the queries align with the system’s knowledge bounds. Such an assessment enables the expert to more accurately determine whether additional retrieval actions are needed. This capability is especially beneficial for complex queries that demand deeper reasoning regarding information sufficiency. The sequential modeling process is formalized as follows:
(8) $\mathcal{Q}_{t}=$ $\mathcal{F}_{QR}(\mathcal{Q}_{t-1},\mathcal{D}^{t}),$
$\mathcal{A}_{t}=$ $\mathcal{F}_{IA}(\mathcal{Q}_{t},\mathcal{D}^{t}).$
Both the retrieval decision-making function $\mathcal{F}_{IA}$ and query reformulation function $\mathcal{F}_{QR}$ are implemented through MLLM generation. To effectively harness the capabilities of MLLMs within these planning sub-tasks, we meticulously engineer tailored prompts.
4.2.2. Retrieval and Generation
At each iterative stage, the selection of retrieval actions determines whether or not and which retrieval API is invoked, either text retrieval or image retrieval. We leverage Google Web Search and Google Image Search as our primary retrieval API service. For each text retrieval request, we retrieve the top-k search results to ensure that only the most relevant information is retained. These results undergo preprocessing through the Jina API framework  [^2] https://github.com/jina-ai/reader, which transforms the raw web content into structured representations better suited for MLLM consumption. The visual retrieval pipeline captures full-page screenshots of search results, employing a set of systematic human-crafted rules processes to eliminate extraneous elements such as white space and original query images. To maintain computational efficiency while preserving result quality, we restrict image retrieval to 3-6 high-confidence results, including their associated contextual captions.
This iterative cycle culminates when the planning expert collectively assesses that the acquired information is sufficiently comprehensive and the formulated query exhibits adequate clarity. Upon reaching this convergence criterion, the planning procedure is finalized, and CogPlanner proceeds to generate the ultimate response.
4.3. Compatibility
The CogPlanner framework is inherently agnostic to the specific model employed, making it easy to be integrated into any MRAG system and immediately enhancing their performance, demonstrating stunning flexibility in real-world applications. The planning expert responsible for query reformulation and determining the appropriate retrieval action can be any MLLM, or even a traditional classification model. For the planning expert, we exclusively employ a diverse set of advanced MLLMs as the foundation. Specifically, we leverage both closed-source APIs and open-source MLLMs. The closed-source models include GPT-4o, while the open-source models consist of the Qwen-VL series (Bai et al., 2023) and the Pixtral series (Dubey et al., 2024).
5. CogBench Construction
In this section, we present CogBench, a benchmark specifically developed for the MRAG planning task. CogBench comprises over 7,000 data samples, with a high-quality test set containing more than 400 samples. This benchmark is designed to facilitate the assessment of the effectiveness of our proposed CogPlanner framework, as well as other MRAG planning frameworks. Moreover, CogBench can be leveraged to enhance the decision-making capabilities of various MLLMs, particularly resource-efficient models, through fine-tuning. In the following subsections, we detail the construction process of CogBench and demonstrate how the benchmark enables lightweight integration of CogPlanner with the Qwen2-VL-7B-Instruct.
5.1. Query Collection
The rapid evolution of MLLMs has created an imperative need to evaluate these systems against increasingly sophisticated user queries that mirror real-world applications. While existing benchmarks (Jiang et al., 2024b; Hu et al., 2024) provide valuable groundwork, we recognized the necessity to extend beyond their scope. To this end, we deliberately incorporate more intricate queries that specifically demand image-based knowledge augmentation. We acquire authentic user interactions through web-sourced screenshots. We first curate a diverse array of topics and structure search queries around these topic words, such as ”CMU posters”. We use Google Image Search to collect image corpus. The resultant image corpus is subject to a stringent filtration process, based on a set of predefined quality criteria and a manual selection process. To generate realistic queries, we leveraged GPT-4o to simulate real users, producing five distinct questions per image that span both factual and open-ended inquiries requiring visual context interpretation. Each query-image pair undergoes manual review by two senior AI research engineers, each bringing at least three years of domain expertise. The engineers select the most compelling 1-5 queries per image that exemplify the strong potential for multimodal retrieval.
5.2. MRAG Planning and Generation
The MRAG planning and generation process is central to the CogPlanner framework, as detailed in Section 4.2. We employ GPT-4o API for executing the planning process. This implementation records each iteration of the planning process, encompassing the series of actions taken, the multimodal document sets retrieved, and the responses generated. Following response generation, the expert annotators conduct a thorough examination to review and regularize the entire chain of actions, and manually annotate the golden answer. Each data sample in the CogBench, therefore, contains a multimodal query, the retrieval actions, reformulated queries at each iterative step, the documents retrieved, and the final golden answer. The CogBench dataset is divided into training and test sets, comprising 5307 and 401 samples, respectively. A particularly noteworthy aspect of the annotation process is the possible variability in the decision chains for identical queries in the test set. While these chains may diverge considerably, the final reformulated queries, derived from the terminal pre-generation iteration, serve as a reflective ”result” of the entire planning procedure. These final queries encapsulate the cumulative informational requirements and explicit intent of the planning process, making it a crucial benchmark for assessing the accuracy of MRAG planning. Thus, the terminal queries are carefully constructed by human annotators.
5.3. CogBench Analysis
As shown in Table 1 and Figure 3, CogBench contains 5718 user queries spanning 9 distinct cognitive domains. We identify several fundamental characteristics that distinguish CogBench from existing benchmarks. (1) Unlike previous benchmarks, which only focus primarily on query-response pairs, CogBench offers a comprehensive record of the entire planning procedure involved in MRAG tasks, thereby facilitating the training of MLLMs. (2) A critical limitation of current MRAG benchmarks lies in their reliance on artificial query construction, primarily through simple entity substitution techniques (Lin and Byrne, 2022; Hu et al., 2024). Such methodologies typically yield responses confined to single entities or numerical values, severely understating the complexity inherent in real-world multimodal interaction scenarios. In contrast, CogBench introduces 24.19% open-ended queries that demand sophisticated, multi-faceted responses encompassing multiple interconnected claims with much longer answer length - 40.13 tokens on average. (3) CogBench incorporates diverse planning procedures that necessitate distinct search strategies at different stages, resulting in varied decision chains across different MLLMs. To quantify this complexity, we also ask the annotators to assess the number of reasoning steps required for resolution (e.g. a single-round of question answering is considered a 1-hop query). Our findings reveal that 79.55% cases explicitly require MRAG Planning, highlighting the sophisticated nature of the reasoning tasks presented in our benchmark.
Table 1. Key statistics of CogBench.
# Query # Domians # Query Len. # Answer Len. # Images
5718 9 8.95 40.13 1381
Reasoning Steps Answer Type
1-hop 2-hop ¿ 2-hop open-ended close-ended
1166 (20.39%) 1882 (32.91%) 2666 (46.62%) 1383 (24.19%) 4334 (75.80%)
Figure 3. Domain distribution of CogBench.
5.4. Lightweight Integration of CogPlanner
To improve the efficiency and reduce the resource requirements of the CogPlanner framework, we aim to achieve a more lightweight integration within MRAG systems. Building upon the CogBench, we introduce a specialized fine-tuning strategy tailored for smaller, resource-efficient MLLMs to broaden the applicability of the CogPlanner framework and mitigate resource constraints. We utilize the Qwen2-VL-7B-Instruct as the backbone and employ the CogBench training set as specialized training data. To maintain a balance between the fine-tuning process and the retention of the Qwen2-VL-7B-Instruct model’s general capabilities, we augment the training dataset with general instruction data at a 1:1 ratio. This ensures that the model benefits from the specialized training required for CogPlanner integration while preserving its broad functionality. The fine-tuned model is referred to as Qwen2-7b-VL-Cog. This fine-tuning methodology is highly adaptable and capable of being applied to any existing MLLM. In our practice, we find that the CogBench fine-tuning process significantly enhances the MRAG planning capabilities of MLLMs, qualified for an effective planning expert. Ultimately, this approach facilitates the development of a lightweight integration of CogPlanner, enhancing MRAG performance while requiring minimal additional computational resources.
Table 2. Performance comparison between CogPlanner and baseline MRAG methodologies on CogBench. Precision and recall are evaluated at the claim level, while the F1 score is assessed at the token level. The diverse planning procedures required by CogBench lead to performance degradation across all fixed pipeline baselines, whereas CogPlanner demonstrates substantial improvements.
Model Reasoning-Steps End-to-End Performance
1-hop 2-hop ¿ 2-hop
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
Origin MLLMs
GPT-4o 0.3349 0.3744 0.1001 0.4438 0.5946 0.3984 0.1617 0.2474 0.1283 0.2907 0.3885 0.2121
Pixtral-Large-Instruct 0.2430 0.4187 0.0519 0.3765 0.5464 0.3451 0.1077 0.2450 0.0713 0.2245 0.3805 0.1581
QVQ-72B-Preview 0.3415 0.2210 0.1056 0.4051 0.3129 0.2148 0.1942 0.1399 0.0637 0.2943 0.2139 0.1224
Qwen2-VL-72B-Instruct 0.3445 0.4885 0.0795 0.3546 0.4494 0.3700 0.1255 0.1700 0.0927 0.2463 0.3278 0.1819
With Fixed Image Retrieval
GPT-4o 0.3173 0.2313 0.1074 0.3675 0.4514 0.3607 0.1788 0.301 0.1113 0.2697 0.3367 0.1932
Pixtral-Large-Instruct 0.2310 0.3927 0.0613 0.2805 0.4451 0.2928 0.1258 0.2491 0.0688 0.1986 0.3435 0.1416
QVQ-72B-Preview 0.1135 0.2313 0.0272 0.1710 0.1977 0.2042 0.0817 0.1024 0.0270 0.1178 0.1603 0.0858
Qwen2-VL-72B-Instruct 0.2630 0.3565 0.0902 0.2692 0.3244 0.2920 0.1645 0.1745 0.0789 0.2193 0.2614 0.1519
With Fixed Text Retrieval
GPT-4o 0.2268 0.1563 0.0572 0.3872 0.3679 0.3708 0.1572 0.1711 0.1049 0.2477 0.2333 0.1833
Pixtral-Large-Instruct 0.0835 0.1422 0.0157 0.1691 0.2173 0.2727 0.0858 0.1301 0.0718 0.1130 0.1615 0.1270
QVQ-72B-Preview 0.2922 0.1854 0.0804 0.2490 0.1662 0.1875 0.0995 0.0808 0.0602 0.1885 0.1305 0.1066
Qwen2-VL-72B-Instruct 0.2056 0.2447 0.0712 0.2666 0.2707 0.3064 0.1163 0.1333 0.0998 0.1844 0.2017 0.1625
CogPlanner with Parallel Modeling
GPT-4o 0.3574 0.3767 0.1046 0.4746 0.5737 0.4283 0.2409 0.2773 0.1365 0.3422 0.3959 0.2267
Pixtral-Large-Instruct 0.2146 0.3708 0.0516 0.4535 0.5229 0.3896 0.2095 0.3034 0.1112 0.2915 0.3900 0.1913
QVQ-72B-Preview 0.3638 0.3644 0.4816 0.4390 0.2960 0.2443 0.2219 0.2522 0.2669 0.3229 0.2897 0.3033
Qwen2-VL-72B-Instruct 0.3550 0.4877 0.0823 0.4588 0.4708 0.3851 0.2164 0.2400 0.1571 0.3245 0.3690 0.2174
CogPlanner with Sequential Modeling
GPT-4o 0.3292 0.3346 0.1014 0.4967 0.5468 0.4326 0.2803 0.3338 0.1526 0.3621 0.4046 0.2349
Pixtral-Large-Instruct 0.2260 0.3639 0.0557 0.3718 0.5432 0.3960 0.0996 0.2996 0.1128 0.2157 0.3936 0.1951
QVQ-72B-Preview 0.3562 0.4595 0.5169 0.4336 0.4407 0.2537 0.1954 0.2118 0.2727 0.3073 0.3384 0.3163
Qwen2-VL-72B-Instruct 0.3688 0.4886 0.0784 0.4294 0.4445 0.3748 0.2157 0.2501 0.1427 0.3179 0.3633 0.2065
6. Experiments
6.1. Experimental Settings
Our experimental evaluation of CogPlanner, conducted on CogBench test set, encompasses two primary dimensions. First, we assess the comprehensive performance of the overall MRAG system to quantify CogPlanner’s effectiveness. Second, we conduct a focused analysis of CogPlanner’s planning procedure to validate our architectural design choices.
6.1.1. End-to-End MRAG Performance
We designed five distinct experimental configurations across all selected MLLMs. The baseline configuration employs the original MLLMs, where multimodal queries are processed directly by the MLLM. We then examine two intermediate configurations: one incorporating fixed image retrieval based on visual query components, and another utilizing fixed text retrieval driven by textual query components. The core evaluation focuses on both parallel and sequential modeling implementations of CogPlanner. For performance metrics, we adopt both token-level and claim-level evaluations, inspired by (Ru et al., 2024). Token-level evaluation is performed using the F1 score, measuring the overlap of common tokens between the generated response and the ground truth. Specifically, we use the NLTK tokenizer to segment the generated answers and ground truth. For claim-level evaluation, we utilize both precision and recall, calculated by first extracting claims from both the golden and generated answers using GPT-4o. Precision measures the proportion of correct claims within the generated responses, while recall evaluates the proportion of correct claims relative to the ground-truth answer claims.
6.1.2. Planning Procedure Performance
In addition to the overall MRAG performance, we examine the efficiency of CogPlanner’s planning procedure, with emphasis on the query reformulation action. We evaluate three distinct approaches: parallel modeling, sequential modeling, and direct reformulating query through GPT-4o with optimized prompt engineering. The comparative analysis employs standard metrics including BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and F1 scores to assess the quality and relevance of the reformulated query outputs against established ground truth.
6.1.3. Implementation Details
We utilize the following backbone MLLMs, GPT-4o, QVQ-72B-Preview (Team, 2024), Qwen2-VL-72B-Instruct (Wang et al., 2024), Pixtral-Large-Instruct (Jiang et al., 2024a), and our fine-tuned Qwen2-7b-VL-Cog. In our retrieval process, we retain the top five results from a web search, each result containing a maximum of 800 tokens. For fine-tuning the Qwen2-7b-VL-Cog model, we leverage the Llama-Factory framework (Zheng et al., 2024) with a learning rate of 2e-6. A cosine learning rate scheduler is employed, and training proceeds for 2 epochs with a batch size of 32 and a warm-up ratio of 0.1. Our experiments are conducted on 8 NVIDIA A800 GPUs.
Table 3. Performance of the query reformulation task across different MLLMs.
Category Model BLEU Rouge F1
Prompting GPT-4o 0.1629 0.4951 0.5375
Parallel GPT-4o 0.1922 0.5266 0.5620
Pixtral-Large-Instruct 0.1678 0.4614 0.5089
QVQ-72B-Preview 0.0702 0.3560 0.3862
Qwen2-VL-72B-Instruct 0.0907 0.4140 0.4472
Sequential GPT-4o 0.1739 0.5050 0.5460
Pixtral-Large-Instruct 0.1773 0.4707 0.5221
QVQ-72B-Preview 0.0532 0.3350 0.3856
Qwen2-VL-72B-Instruct 0.0918 0.4266 0.4643
6.2. CogPlanner Performance
6.2.1. End-to-end Performance
Table 2 presents a comprehensive comparison of the end-to-end performance comparison of various MLLMs integrated with current MRAG methodologies and our proposed CogPlanner. The following observations can be drawn based on these results: (1) Enhanced Performance with CogPlanner: Notably, CogPlanner with GPT-4o consistently yields best performance compared to all other configurations. Specifically, CogPlanner results in a significant improvement over MRAG baselines, with performance gains ranging from 12.4% to 52.5%, as measured by the average end-to-end performance metrics. This enhancement is attributed to its ability to decompose and refine complex, multimodal queries. By simplifying these queries, CogPlanner facilitates the dynamic determination of necessary retrieval actions, thereby ensuring the acquisition of accurate, complementary information. These results highlight the critical role of MRAG Planning in optimizing the performance of MRAG systems. (2) Weakness of Fixed Search Strategies: The two fixed search strategies, while offering some improvements in specific metrics, generally exhibit a negative impact when compared to direct responses generated by MLLMs. As anticipated, these rigid search actions, particularly when applied to concise queries and screenshot images within multimodal queries, introduce noise that misleads downstream MLLMs rather than providing useful information. The noisy retrieval results tend to obscure the relevant information, thereby diminishing the effectiveness of the system. These findings further validate the motivation behind the development of the CogPlanner framework. (3) Comparison of Parallel and Sequential Modeling Approaches: Comparing our parallel and sequential modeling methodologies, they show comparable performance, each demonstrating significant improvements over traditional MRAG systems. However, the sequential modeling approach does not exhibit substantial advantage over the parallel approach. This suggests that the current capabilities of MLLMs are insufficient for accurately evaluating complementary information in a sequential manner (4) Benefits on Multihop Query: The performance of CogPlanner shows variability across 1-3 hop query categories, with more pronounced improvements observed in multi-hop queries. This pattern suggests that the iterative and adaptive planning process of CogPlanner is particularly beneficial in scenarios that require the retrieval of information from multiple sources and involve multi-step reasoning to formulate a complete and accurate response.
Table 4. End-to-end performance comparison of Qwen2-VL-72B-Instruct and Qwen2-7b-VL-Cog as planning experts, with response generation models consistently using Qwen2-VL-72B-Instruct.
Planning Expert Precision Recall F1 Avg
CogPlanner With Parallel Modeling
Qwen2-VL-72B-Instruct 0.3245 0.3690 0.2174 0.3036
Qwen2-7b-VL-Cog 0.3197 0.3265 0.2146 0.2869
CogPlanner With Sequential Modeling
Qwen2-VL-72B-Instruct 0.3179 0.3633 0.2065 0.2959
Qwen2-7b-VL-Cog 0.3250 0.3310 0.2138 0.2900
6.2.2. Query Reformulation Performance
As illustrated in Table 3, we evaluate the query reformulation performance of CogPlanner in comparison with various backbones, including direct prompt engineering approaches with these MLLMs. Among the configurations tested, GPT-4o with parallel modeling emerges as the highest-performing setting. Notably, CogPlanner demonstrates its ability to significantly enhance the accuracy and informativeness of the original multimodal queries, thereby improving the overall query formulation process. This improvement underscores the core objective of query reformulation within the context of MRAG Planning, highlighting the robustness and effectiveness of the CogPlanner framework. Moreover, the results reveal that the query reformulation performance is highly contingent upon the selection of the planning expert MLLMs.
Table 5. The efficiency evaluation of CogPlanner overhead in MRAG systems, with the ratio calculated based on the average token count of the generated response for each experimental setting.
Planning Expert Category # Para Tokens # Total Tokens
Qwen2-VL-72B-Instruct Parallel 8.90 (13.7%) 9.76 (14.9%)
Sequential 13.56 (21.6%) 13.56 (21.6%)
Qwen2-7b-VL-Cog Parallel 6.59 (8.5%) 7.58 (9.8%)
Sequential 7.59 (11.9%) 7.59 (11.9%)
6.3. Lightweight Model Performance
In this section, we assess the performance of our fine-tuned Qwen2-7b-VL-Cog from both effectiveness and efficiency perspectives.
6.3.1. Model Evaluation
To ensure a fair comparison, we use the Qwen2-VL-72B-Instruct as the reference model for final answer generation. Both Qwen2-VL-72B-Instruct and our fine-tuned Qwen2-7b-VL-Cog are utilized as planning experts within the CogPlanner framework, facilitating an assessment of whether smaller MLLMs, specifically through our tailored supervised fine-tuning procedure, can effectively manage the complexities inherent in MRAG planning tasks. The results, as presented in Table 4, underscore the critical role of model selection in determining the overall performance. While larger MLLMs typically demonstrate superior results, our Qwen2-7b-VL-Cog closely approximates the performance of Qwen2-VL-72B-Instruct across most evaluation metrics. This finding serves to validate the efficacy of our fine-tuning strategy. Under the CogPlanner framework, our results demonstrate that it is indeed possible to deploy a resource-efficient planning expert and lead to enhanced performance of the MRAG system. Specifically, the Qwen2-7b-VL-Cog emerges as a compelling alternative.
6.3.2. Efficiency Evaluation
We conduct a comparison of the efficiency of CogPlanner between Qwen2-VL-72B-Instruct and Qwen2-7b-VL-Cog within the context of MRAG systems, focusing on the overall cost of the planning procedure. We utilize token generation as the primary metric for evaluating efficiency. Specifically, the ”Total Tokens” refers to the cumulative number of tokens generated by the MLLMs during the entire planning process. Additionally, we introduce the ”Para Token” metric as a near-latency measure. Given that the parallel modeling paradigm uses two threads for decision-making across two actions, we consider the maximum token count generated during each round of planning as the cost for that round. To provide a more comprehensive understanding of the additional computational cost incurred by incorporating CogPlanner, we calculate the average number of tokens in the generated response and compute the ratio between the tokens used in planning versus those needed for answering. The evaluation results are summarized in Table 5. Our findings indicate that Qwen2-7b-VL-Cog is a significantly more lightweight module for MRAG systems compared to Qwen2-VL-72B-Instruct, introducing only a 10% increase in computational cost. This result underscores the minimal additional overhead imposed by the lightweight integration of CogPlanner into MRAG systems, highlighting its suitability for real-world applications. Furthermore, the parallel execution model exhibits superior performance relative to the sequential modeling approach across both efficiency metrics, aligning with the design expectations.
6.4. Analysis
Table 6. The proportion of retrieval actions of different methodologies, # No, # Text, # Image represents no search, text search and image search, respectively.
Model Category # No # Text # Image
Pixtral-Large-Instruct Parallel 11.51% 65.24% 23.25%
Sequential 8.43% 84.28% 18.00%
Qwen2-VL-72B-Instruct Parallel 13.38% 59.87% 26.75%
Sequential 11.46% 66.07% 22.47%
Qwen2-7b-VL-Cog Parallel 5.25% 80.43% 14.32%
Sequential 5.25% 80.42% 14.43%
Figure 4. The average number of retrieval actions of different MLLMs across queries with different reasoning steps.
6.4.1. Analysis on planning procedure of Cogplanner
We conduct an analysis to explore the adaptive decision-making capabilities of CogPlanner, which emulates human cognition by tailoring its planning processes to the specific knowledge bases of different MLLMs. Specifically, we examine the length of decision chains and the distribution of retrieval actions across the Pixtral-Large-Instruct, Qwen2-VL-72B-Instruct, and Qwen2-7b-VL-Cog. The results are summarized in Table 6 and Figure 4. The following observations can be drawn from these analyses: (1) As shown in Figure 4, the distribution of retrieval actions indicates that all the MLLMs tend to perform more actions than are typically expected by human annotators. The expected behavior would be a gradual progression of actions. However, MLLMs generally opt to acquire more information and make conservative decisions in an attempt to ensure accuracy (Chen et al., 2024; Zhang et al., 2024a). Among the models examined, the Qwen2-VL-72B-Instruct exhibits the most pronounced mismatch, performing over two rounds of processing even for 1-hop queries, which is counterintuitive when compared to its behavior on more complex, multi-hop queries. In contrast, the Qwen2-7b-VL-Cog model produces the most reasonable number of retrieval actions. Additionally, comparing sequential and parallel modeling paradigms reveals a substantial reduction in redundant retrieval actions. This is attributed to the query reformulation step, which allows the CogPlanner to better evaluate whether further search is genuinely necessary, resembling a reflective thought process. (2) From the data presented in Table 6, it is evident that the Qwen2-7b-VL-Cog model predominantly relies on textual search. This trend is also observable when comparing the two modeling paradigms. These results suggest that text-based retrieval remains the most preferred and fundamental method for acquiring information in MLLMs. Furthermore, the prevalence of redundant retrieval actions in this domain could be attributed to the tendency of models to perform additional web searches for reassurance, which does not necessarily harm overall performance. (3) Notably, most tasks appear to be resolved around 2-hop retrieval steps, even for more complex queries requiring greater than 2 hops. Determining an optimal retrieval strategy that aligns with the model’s knowledge base and reasoning capabilities remains a challenging task for current MLLMs.
7. Conclusion
This work introduces Multimodal Retrieval Augmented Generation Planning (MRAG Planning) to address the limitations of current MRAG frameworks. Our research underscores the importance of dynamically optimizing the information acquisition and query reformulation processes. The CogPlanner framework leverages decision-making to refine queries and select appropriate retrieval strategies, minimizing redundant retrieval and enhancing response quality. CogPlanner offers flexible modeling approaches and integrates seamlessly with existing MRAG systems. Additionally, we propose the CogBench benchmark to assess MRAG Planning’s decision-making capabilities, filling a gap in current evaluation methods. The experimental results validate the efficacy of MRAG Planning and CogPlanner, showcasing substantial improvements in performance with minimal additional computational costs. This work paves the way for more adaptive and effective MRAG frameworks.